Find the x and y intercepts and graph 7x-9y=63 by plotting the intercepts. Please show all of your work.
Answer by mananth(16949)

7x-9y=63
9y=7x-63
y=7x/9-7
x intercept, y=0
7x-0=63
x=9 , (9,0)
--
Y intercept x =0
-9y=63
y=-7, (0,-7)